find the value of k such the graph of 3x + k = 5y has an x-intercept of (-4,0)
Answer by stanbon(75887)

find the value of k such the graph of 3x + k = 5y has an x-intercept of (-4,0)
---
Substitute for x and y and solve for "k".
---
3*-4 + k = 5*0
-12 + k = 0
k = 12
--------------------
Cheers,
Stan H.